The horse is tethered to a rope, at one end of a square corral(outside the corral) 10 feet on a side. The horse can graze at a distance of 18feet from the corner of the corral where the rope is tied. What is the total grazing area of the horse.
Answer by checkley77(12844)

SETTING UP THIS PROBREM WITH 4 PARTIAL CIRCLES.
ONE A SEMICIRCLE OF RADIUS 18 FT.
ONE QUARTER CIURCLE OF RADIUS 18 FT.
AND 2 QUARTER CIRCLES OF RADIUS 8 FT.
AREA=PI*R^2
.5*3.14*18^2=508.68 FT^2
.25*3.14*18^2=254.34 FT^2
2*.25*3.14*8^2=100.48 FT^2
TOTAL AREA IS:
508.68+254.34+100.48=863.5 FT^2 ANSWER.